### v1

# Theoretical Limits of Protocols for Distinguishing Different Unravelings

Abstract
The evolution of an open quantum system is often described by a master equation, which governs the dynamics of the statistical operator. The dynamics of a master equation can be expressed as an ensemble average over stochastic trajectories, known as stochastic unravelings, which correspond to different measurement schemes and provide a wavefunction-level description of the system’s evolution. The fact that a given master equation admits multiple unravelings raises the question of whether these different stochastic descriptions can be operationally distinguished, as recently suggested. We analyze this possibility: we show that while unraveling-dependent quantities indeed differ at the mathematical level, as already known in the literature, they can only be computed once the measurement procedure implementing the unraveling is already given, rendering them inaccessible when the procedure is unknown.
I Introduction
In quantum mechanics, especially in the theory of open quantum systems Breuer and Petruccione (2002); Rivas and Huelga (2012), the information about a system of interest is encoded in the statistical operator $\hat{\rho}_{t}$. The seminal works of V. Gorini, A. Kossakowski, G. Sudarshan Gorini et al. (1976), and G. Lindblad Lindblad (1976) identified, under general assumptions, its dynamical evolution; limiting to a single Hermitian Lindblad operator $\hat{L}$, the so-called GKLS master equation reads:
$\frac{\text{d}\hat{\rho}_{t}}{\text{d}t}=-\frac{i}{\hbar}[\hat{H},\hat{\rho}_{t}]-\frac{\lambda}{2}[\hat{L},[\hat{L},\hat{\rho}_{t}]],$ (1)
where $\hat{H}$ is the Hamiltonian of the system and $\lambda$ a positive coupling constant.
Over the years, interest has increased in stochastic unravelings of the master equation (1); these are suitably modified Schrödinger’s equations with the addition of stochastic terms, whose random solutions $|\psi_{\omega,t}\rangle$, often called quantum trajectories Carmichael (2009), are such that the conditional state Cao et al. (2019) $\hat{\rho}_{\omega,t}=\ket{\psi_{\omega,t}}\bra{\psi_{\omega,t}}$ | when averaged with respect to the noise, is a solution of the master equation above, i.e. ${\mathbb{E}}_{\omega}[\hat{\rho}_{\omega,t}]=\hat{\rho}_{t}$, where ${\mathbb{E}_{\omega}}[\cdot]$ denotes the stochastic average. As a notational remark, in the following we will term conditional any quantity, such as the state or the quantum expectation value of an observable, before taking the stochastic average; otherwise that quantity is meant to be averaged over the noise. Moreover, we will omit to indicate explicitly the dependence of a conditional quantity on the noise ($\omega$), unless confusion may arise.
Stochastic Schrödinger equations (SSEs) appear in a variety of contexts Wiseman and Milburn (2009, 1993); Carmichael (2009); Dalibard et al. (1992); Nielsen and Chuang (2010); Preskill (2018); Breuer and Petruccione (2002); Gisin and Percival (1992); Mølmer et al. (1993); Daley (2014); Diósi (1998); Ghirardi et al. (1990); Kafri et al. (2014); Tilloy and Diósi (2016, 2017); Weiss (2012); Hänggi et al. (1990); Wiseman and Diósi (2001); Caiaffa et al. (2017); Jacobs (2014); Wiseman (1996); Carmichael (2007); Piccitto et al. (2022); Cao et al. (2019); Vovk and Pichler (2024); Piccitto et al. (2024); Bassi and Ippoliti (2006); Genoni et al. (2016); Wu et al. (2024); Piñol et al. (2024); Albarelli et al. (2017); Magrini et al. (2021); Zhang and Mølmer (2017); Belenchia et al. (2020); Zhang et al. (2017). In quantum optics, they are used to describe the dynamics of systems under continuous measurement Wiseman and Milburn (2009, 1993), for different photon detection schemes Carmichael (2009); Dalibard et al. (1992). In quantum computation, they are employed to analyze the impact of noise and errors on qubits, in light of fault-tolerant quantum computing Nielsen and Chuang (2010); Preskill (2018). Within the theory of open quantum systems, they allow to describe the reduced system dynamics at the wave function level Breuer and Petruccione (2002); Gisin and Percival (1992). Numerical simulations based on SSEs, including quantum Monte Carlo methods, enable the simulation of larger quantum systems, making SSEs a common tool for tackling computationally intensive problems Mølmer et al. (1993); Daley (2014). Stochastic Schrödinger equations are also used in the foundations of quantum mechanics to model the spontaneous collapse of the wave function Diósi (1998); Ghirardi et al. (1990); more recently, they have been use to model the coupling between classical gravity and quantum matter Kafri et al. (2014); Tilloy and Diósi (2016, 2017) offering an alternative to quantum gravity.
The association between master equations and stochastic unravelings is not unique: in general, there are infinite ways to unravel a master equation via SSEs Wiseman and Diósi (2001); Caiaffa et al. (2017), as also shown later on. For example, in quantum optics, the detection of photons at random times leaking out of an optical cavity is described by a stochastic equation for the state of the system, with the randomness encoded in a Poisson increment Jacobs (2010, 2014). In homodyne detection, a photodetector is used to perform a Gaussian continuous measurement of an optical mode, where instead of a direct detection, light is first interfered (with light of the same frequency) in a beam splitter. The resulting homodyne stochastic equation is defined in terms of a Wiener process Wiseman (1996); Carmichael (2007); Jacobs (2014). These are two different measurement processes; yet at the statistical level, when averaged over the noise, they are indistinguishable from each other Jacobs (2014).
Within the context of stochastic dynamics, quantities such as the mean value or the variance of observables clearly play a crucial role. In the case of SSEs, there are two types of such quantities: those such as $({\mathbb{E}}_{\omega}[\langle\hat{O}\rangle_{t}])^{n}\equiv(\text{Tr}[\hat{O}\hat{\rho}_{t}])^{n}$ (since $\hat{\rho}_{t}={\mathbb{E}}_{\omega}[|\psi_{t}\rangle\langle\psi_{t}|]$), which by construction can be read directly from the density matrix without the need of unraveling, except possibly for numerical reasons, as mentioned before; and those like ${\mathbb{E}}_{\omega}[\langle\hat{O}\rangle_{t}^{n}]$ ($n>1$) which depend nonlinearly on the conditional state of the system $\hat{\rho}_{\omega,t}$ so they cannot be expressed in terms of the density matrix. The latter quantities crucially depend on the chosen unraveling. The typical example of this second type of quantities is ${\mathbb{E}}_{\omega}[\Sigma_{t}(\hat{O}_{i},\hat{O}_{j})]$, with $\Sigma_{t}(\hat{O}_{i},\hat{O}_{j})=\frac{1}{2}\langle\{\hat{O}_{i},\hat{O}_{j}\}\rangle_{t}-\langle\hat{O}_{i}\rangle_{t}\langle\hat{O}_{j}\rangle_{t}$, which for example appear in quantum optics Genoni et al. (2016); Wu et al. (2024); Piñol et al. (2024); Albarelli et al. (2017); Magrini et al. (2021); Zhang and Mølmer (2017); Belenchia et al. (2020); Zhang et al. (2017) in the form of the quantum Riccati equation  [^1] See Appendix C where we derive the dynamics for a monitored particle in a harmonic potential in the different unravelings for the covariance matrix for the quadratures of a mechanical system in the limit of Gaussian states.
Recently, it has been suggested Piñol et al. (2024) that these nonlinear quantities can be used to operationally distinguish different unravelings of a given master equation. This question is both interesting and relevant to the theory of open quantum systems. However, here we show that this is not the case.
The reason is that nonlinear quantities are either reconstructed from the unraveling associated with the measurement procedure, and thus quantify differences that are embodied in the choice of the unraveling—i.e., the operational distinction is already made in selecting the measurement procedure, not in computing the associated nonlinear quantities—or they are not accessible at all. The latter point is particularly crucial, as their accessibility would enable the development of a protocol for superluminal signaling, which is (supposed to be) forbidden.
II Nonlinear and linear unravelings of the master equation
There exists a large variety of unravelings of a master equation such as Eq. (1); for the purposes of our work, in the following we will consider the following family of continuous unravelings Adler and Bassi (2007):
$\begin{split}\text{d}\ket{\psi_{t}}&\!=\!\left(\!-\frac{i}{\hbar}\hat{H}dt\!-\!\frac{\lambda}{2}(|\xi|^{2}\hat{L}^{2}-2\xi\xi^{\text{\tiny{R}}}\hat{L}\langle\hat{L}\rangle_{t}+(\xi^{\text{\tiny{R}}})^{2}\langle\hat{L}\rangle_{t}^{2})\text{d}t\right.\\ &\left.+\sqrt{\lambda}(\xi\hat{L}-\xi^{\text{\tiny{R}}}\langle\hat{L}\rangle_{t})\text{d}W_{t}\right)\ket{\psi_{t}}\end{split}$ (2)
where $\langle\hat{L}\rangle_{t}=\bra{\psi_{t}}\hat{L}\ket{\psi_{t}}$, $W_{t}$ is a standard Wiener process and $\xi=\xi^{\text{\tiny{R}}}+i\xi^{\text{\tiny{I}}}$ is a complex parameter such that $\xi^{\text{\tiny{R}}}\geq 0$ and $|\xi|=1$. Throughout this work we will use the Itô formalism. With the exception of the limiting case $\xi^{\text{\tiny R}}=0$, one can show  [^2] See Appendix A for the details that the above family of unravelings corresponds to different types of Gaussian continuous measurement Jacobs (2014), with measurement record
$\text{d}y_{t}=\xi^{\text{\tiny{R}}}\langle\hat{L}\rangle_{t}\text{d}t+\frac{1}{2\sqrt{\gamma}}\text{d}W_{t}.$ (3)
From a known initial state, using Eqs. (3) and (2) one can always reconstruct the evolution of the system: at a given time $t$ one can compute the conditional expectation $\langle\hat{L}\rangle_{t}$ through the conditional state $|\psi_{t}\rangle$; then, by measuring the signal $y_{t}$, Eq. (3) allows to reconstruct the noise contribution $\text{d}W_{t}$, which can be used to fully determine the state at a time $t+\text{d}t$ via Eq. (2).
These unravelings lead to different evolutions for the conditional wave function. To see this, in what follows we will consider two extreme cases of Eq. (2): $\xi=\xi^{\text{\tiny{R}}}$ (i.e. $\xi^{\text{\tiny{I}}}=0$), corresponding to a standard continuous quantum measurement; and $\xi=\xi^{\text{\tiny{I}}}$ (i.e. $\xi^{\text{\tiny{R}}}=0$), corresponding to a linear dynamics with the addition of a stochastic potential.
The first choice, as mentioned above, arises in the context of Gaussian continuous measurements Jacobs (2014); Wiseman and Milburn (2009), as well as dynamical collapse models Bassi and Ghirardi (2003); Bassi et al. (2013). The dynamics is nonlinear in the wavefunction, but norm-preserving:
$\text{d}\ket{\psi^{\text{\tiny A}}_{t}}=\left(-\frac{i}{\hbar}\hat{H}\text{d}t-\frac{\lambda}{2}(\hat{L}^{\text{\tiny A}}_{t})^{2}\text{d}t+\sqrt{\lambda}\hat{L}^{\text{\tiny A}}_{t}\text{d}W_{t}\right)\ket{\psi^{\text{\tiny A}}_{t}},$ (4)
where $\hat{L}^{\text{\tiny A}}_{t}\equiv\hat{L}-\langle\hat{L}\rangle^{\text{\tiny A}}_{t}$. The second and third terms induce the localization of the wave function towards an eigenstate of the operator $\hat{L}$, which occurs randomly in accordance with the Born probability rule.
The second choice, in contrast to the previous one, leads to a linear dynamics:
$\text{d}\ket{\psi^{\text{\tiny B}}_{t}}=\left(-\frac{i}{\hbar}\hat{H}\text{d}t-\frac{\lambda}{2}\hat{L}^{2}\text{d}t-i\sqrt{\lambda}\hat{L}\text{d}W_{t}\right)\ket{\psi^{\text{\tiny B}}_{t}}.$ (5)
The second term, proportional to $\hat{L}^{2}$, is the Itô correction to the otherwise unitary evolution encoded both in the deterministic contribution of the Hamiltonian $\hat{H}$, as well as the stochastic term proportional to $\hat{L}$, and disappears from the solution of the equation. Also this unraveling is norm-preserving, however, it does not lead to the collapse of the wave function.
Let us now consider an observable $\hat{O}$ of the system. By construction, the dynamical evolution of the conditional expectations $\langle\hat{O}\rangle^{\text{\tiny A}}_{t}$ and $\langle\hat{O}\rangle^{\text{\tiny B}}_{t}$ differs in the two cases  [^3] For the unraveling in Eq. (4), make the substitution $\xi=1$ in Eq. (33). For the unraveling in Eq. (5), make the substitutions $\xi=-i$ in Eq. (33). but coincides once averaged over all realizations of the noise, since $\mathbb{E}_{\omega}[\langle\hat{O}\rangle^{\text{\tiny A}}_{t}]=\text{Tr}[\hat{O}\hat{\rho}_{t}]=\mathbb{E}_{\omega}[\langle\hat{O}\rangle^{\text{\tiny B}}_{t}]$.
As mentioned before, one may distinguish between different unravelings through the examination of quantities that cannot be expressed in terms of the density operator $\hat{\rho}_{t}$. An example of this is $\langle\hat{O}\rangle^{2}_{t}$. For the unraveling in Eq. (4) we have:
$\begin{split}\text{d}{\langle\hat{O}\rangle_{t}^{\text{\tiny A}}}^{2}&=2\langle\hat{O}\rangle_{t}^{\text{\tiny A}}\left(-\frac{i}{\hbar}\langle[\hat{O},\hat{H}]\rangle_{t}^{\text{\tiny A}}-\frac{\lambda}{2}\langle[\hat{L},[\hat{L},\hat{O}]]\rangle_{t}^{\text{\tiny A}}\right)\text{d}t\\ &+\lambda(\langle\{\hat{O},\hat{L}\}\rangle_{t}^{\text{\tiny A}}-2\langle\hat{O}\rangle_{t}^{\text{\tiny A}}\langle\hat{L}\rangle_{t}^{\text{\tiny A}})^{2}\text{d}t\\ &+2\langle\hat{O}\rangle_{t}^{\text{\tiny A}}\sqrt{\lambda}(\langle\{\hat{O},\hat{L}\}\rangle_{t}^{\text{\tiny A}}-2\langle\hat{O}\rangle_{t}^{\text{\tiny A}}\langle\hat{L}\rangle_{t}^{\text{\tiny A}})\text{d}W_{t}\end{split}$ (6)
whereas the unraveling in Eq. (5) leads to the following equation:
$\begin{split}\text{d}{\langle\hat{O}\rangle_{t}^{\text{\tiny B}}}^{2}&=2\langle\hat{O}\rangle_{t}^{\text{\tiny B}}\left(-\frac{i}{\hbar}\langle[\hat{O},\hat{H}]\rangle_{t}^{\text{\tiny B}}\text{d}t-\frac{\lambda}{2}\langle[\hat{L},[\hat{L},\hat{O}]]\rangle_{t}^{\text{\tiny B}}\right)\text{d}t\\ &-\lambda{\langle[\hat{L},\hat{O}]\rangle_{t}^{\text{\tiny B}}}^{2}\text{d}t+2i\langle\hat{O}\rangle_{t}^{\text{\tiny B}}\sqrt{\lambda}\langle[\hat{L},\hat{O}]\rangle_{t}^{\text{\tiny B}}\text{d}W_{t}.\end{split}$ (7)
In taking the average over the noise, due to the properties of the Wiener process, the terms proportional to $\text{d}W_{t}$ vanish. What remains is different in the two cases, as we will explicitly show with the following two examples.
II.1 Free-particle
As a first example, we consider a free particle of mass $m$ undergoing decoherence in position; the simplest choice for Eq. (1) is to take
$\hat{H}=\frac{\hat{p}^{2}}{2m},\qquad\hat{L}=\hat{x}.$ (8)
We will be interested, for each of the two unravelings, in the time evolution of the spread in position: $\mathbb{E}_{\omega}[\Sigma_{t}(\hat{x})]$ with $\Sigma_{t}(\hat{x})~{}=~{}\langle\hat{x}^{2}\rangle_{t}-\langle\hat{x}\rangle^{2}_{t}$. Let us restrict the analysis to Gaussian states; it is not difficult to show that:
${\psi}_{t}(x)=\exp\left[-a_{t}(x-\overline{x}_{t})^{2}+i\overline{k}_{t}x+\gamma_{t}\right]$ (9)
is solution for both unravelings, where $a_{t}$ and $\gamma_{t}$ are complex functions of time, whereas $\overline{x}_{t}$ and $\overline{k}_{t}$ are real. It can be shown  [^4] See Appendix B, where we present the explicit time dependence of $a_{t}$, $\gamma_{t}$, $\overline{x}_{t}$, $\overline{k}_{t}$, and we derive the free particle dynamics in the different unravelings that for the nonlinear unraveling of Eq. (4), with the choices in Eq. (8) for $\hat{H}$ and $\hat{L}$, we have:
$\Sigma^{\text{\tiny A}}_{t}(\hat{x})=\frac{1}{4}\frac{\cosh[2(b^{\text{\tiny{R}}}t+k^{\text{\tiny{R}}})]+\cos[2(b^{\text{\tiny{I}}}t+k^{\text{\tiny{I}}})]}{c^{\text{\tiny{R}}}\sinh[2(b^{\text{\tiny{R}}}t+k^{\text{\tiny{R}}})]-c^{\text{\tiny{I}}}\sin[2(b^{\text{\tiny{I}}}t+k^{\text{\tiny{I}}})]},$ (10)
with $b,c$ and $k$ constants 555cf. Eq. (38). On the other hand, for the linear unraveling, we obtain:
$\Sigma^{\text{\tiny B}}_{t}(\hat{x})=\frac{1}{4}\frac{(m-2\hbar ta_{0}^{\text{\tiny{I}}})^{2}+4\hbar^{2}t^{2}(a_{0}^{\text{\tiny{R}}})^{2}}{m^{2}a_{0}^{\text{\tiny{R}}}}.$ (11)
Note that in both cases $\Sigma_{t}(\hat{x})$ does not depend on the noise, therefore the stochastic average is ineffective.
Both quantities differ from the variance $\text{Var}_{t}(\hat{x})=\mathbb{E}_{\omega}[\langle\hat{x}^{2}\rangle_{t}]-(\mathbb{E}_{\omega}[\langle\hat{x}\rangle])^{2}_{t}=\text{Tr}[\hat{x}^{2}\hat{\rho}_{t}]-(\text{Tr}[\hat{x}\hat{\rho}_{t}])^{2}$  [^6] In general, we have $\mathbb{E}_{\omega}[\text{Var}(\hat{O}_{i},\hat{O}_{j})]=\frac{1}{2}\mathbb{E}_{\omega}\langle\{\hat{O}_{i},\hat{O}_{j}\}\rangle_{t}-\mathbb{E}_{\omega}[\langle\hat{O}_{i}\rangle_{t}]\mathbb{E}_{\omega}[\langle\hat{O}_{j}\rangle_{t}]$, obtained from the density matrix, which by construction is unraveling-independent. Straightforward calculations show that
$\text{Var}_{t}(\hat{x})$ $=\frac{(m-2\hbar ta_{0}^{\text{\tiny{I}}})^{2}+4\hbar^{2}t^{2}(a_{0}^{\text{\tiny{R}}})^{2}}{4m^{2}a_{0}^{\text{\tiny{R}}}}+\lambda\frac{\hbar^{2}}{m^{2}}\frac{t^{3}}{3},$ (12)
which matches the the well-known expression in the decoherence literature Joos et al. (2013) (for $a_{0}^{\text{\tiny{I}}}=0$).
Fig. 1 shows the time evolution of time of $\Sigma^{\text{\tiny A}}_{t}(\hat{x})$, $\Sigma^{\text{\tiny B}}_{t}(\hat{x})$, as well as of $\text{Var}_{t}(\hat{x})$, for given initial values of the parameters. For small times, they roughly coincide, because the free part of the dynamics, which is the same in all cases, dominates over the stochastic part. For larger times, the three quantities depart. $\Sigma^{\text{\tiny A}}_{t}(\hat{x})$ settles to a finite value: the nonlinear unraveling collapses the wave function, counteracting the free expansion, and therefore the Gaussian state reaches asymptotically a finite spread. The linear unraveling does not induces the collapse of the wave function, but acts as a random position-dependent potential, therefore the spread $\Sigma^{\text{\tiny B}}_{t}(\hat{x})$ keeps increasing over time. The same occurs for $\text{Var}_{t}(\hat{x})$ but at a faster rate.
Figure 1: Time evolution of the unraveling-independent variance $\text{Var}_{t}(\hat{x})$ (Eq. (12), green line), and of $\Sigma_{t}(\hat{x})$ for the nonlinear unraveling ($\Sigma_{t}^{\text{\tiny{A}}}(\hat{x})$ in Eq. (10), blue line) and the linear unraveling ($\Sigma_{t}^{\text{\tiny{B}}}(\hat{x})$ in Eq. (11), yellow line). For illustrative purposes, we use the parameters: $\overline{x}_{0}~{}=~{}0$, $\overline{k}_{0}~{}=~{}0$, $m~{}=~{}10^{-15}\,\mathrm{kg}$, $a_{0}~{}=~{}0.25\times 10^{9}\mathrm{m^{-2}}$ (therefore $\Sigma^{\text{\tiny A}}_{0}(\hat{x})=\Sigma^{\text{\tiny B}}_{0}(\hat{x})=\text{Var}_{0}(\hat{x})=10^{-9}\mathrm{m^{2}}$), $\lambda~{}=~{}10^{23}\mathrm{m^{-2}\,Hz}$. The three quantities evolve differently over time.
II.2 Spin-$1/2$ particle
Let us now consider a spin-$1/2$ system; the dynamical evolution is still given again by Eq. (1), where now we set:
$\hat{H}=\hbar\nu\hat{\sigma}_{z},\qquad\hat{L}=\hat{\sigma}_{z}$ (13)
with $\hat{\sigma}_{z}$ the Pauli matrix and $\nu$ a coupling frequency. In this case, the solution of the stochastic equations for both the nonlinear and linear unraveling can be written explicitly. For the nonlinear unraveling given by Eq. (4), the solution is  [^7] See Appendix D, where we derive an explicit solution for the Spin-$1/2$ particle in the nonlinear unraveling:
|$\ket{\psi^{\text{\tiny A}}_{t}}\!\!=\!\!\frac{\exp\!\left[\left(-i\nu t+2\lambda\int_{0}^{t}{\langle\hat{\sigma}_{z}\rangle}_{t}^{\text{\tiny A}}\text{d}s+\sqrt{\lambda}W(t)\right)\hat{\sigma}_{z}\right]\!\ket{\psi_{0}}}{\bra{\psi_{0}}\exp[(2\sqrt{\lambda}W(t)+4\lambda\int_{0}^{t}\langle\hat{\sigma}_{z}\rangle_{t}^{\text{\tiny A}}\text{d}s)\hat{\sigma}_{z}]\ket{\psi_{0}}^{1/2}},$ (14)
where the non-Hermitian part of the dynamics randomly drives the initial state towards one of the two eigenstates of $\hat{\sigma}_{z}$, with a probability given by the Born rule; this is shown in Fig. 2.
Figure 2: Time evolution of $\langle\hat{\sigma}_{z}\rangle^{\text{\tiny{A}}}_{t}$ for ten different realizations of the noise, given the initial state |$\ket{\psi_{0}}=\frac{1}{2}\ket{\uparrow}+\frac{\sqrt{3}}{2}\ket{\downarrow}$. After an initial fluctuating motion, each trajectory converges either to the positive or the negative eigenstate of $\hat{\sigma}_{z}$. In the present case, four trajectories converge to the positive eigenstate, and six to the negative one, in good accordance with the Born rule. All constants ($\hbar,\nu,\lambda$) have been set equal to 1.
On the more formal level, following the same argument as in in Ref. Adler et al. (2001) one can show that  [^8] See Appendix  E, where we show the collapse of the wavefunction in the nonlinear unraveling:
$\lim_{t\rightarrow\infty}\Sigma^{\text{\tiny A}}_{t}(\hat{\sigma}_{z})=\lim_{t\rightarrow\infty}[\langle\hat{\sigma}_{z}^{2}\rangle^{\text{\tiny A}}_{t}-{\langle\hat{\sigma}_{z}\rangle^{\text{\tiny A}}_{t}}^{2}]=0,$ (15)
for any initial state. The conditional spread of $\hat{\sigma}_{z}$ vanishes over time, for almost every realization of the noise, meaning that the state converges towards one of its eigenstates.
In contrast, for the linear unraveling given by Eq. (5), the solution is Jacobs (2010):
|$\ket{\psi^{\text{\tiny B}}_{t}}=\exp\left[\left(-i\nu t-i\sqrt{\lambda}W(t)\right)\hat{\sigma}_{z}\right]\ket{\psi_{0}};$ (16)
now the evolution is unitary and there is no collapse of the wave function. On top of the deterministic rotation controlled by the Hamiltonian $\hat{H}$, the components of the vector, with respect to the eigenstates of $\hat{\sigma}_{z}$, receive random relative phases, which cause decoherence in the $\hat{\sigma}_{z}$ basis. In fact, now the conditional variance $\Sigma^{\text{\tiny B}}_{t}(\hat{\sigma}_{z})$ is constant over time:
$\Sigma^{\text{\tiny B}}_{t}(\hat{\sigma}_{z})=\Sigma^{\text{\tiny B}}_{0}(\hat{\sigma}_{z})=1-\langle\psi_{0}|\hat{\sigma}_{z}|\psi_{0}\rangle^{2};$ (17)
being a deterministic quantity, the average over the noise is ineffective.
We see once again that the two nonlinear variances, referring to the two unravelings, are radically different. In this case, the variance $\text{Var}_{t}(\hat{\sigma}_{z})$, which can be computed from the Lindblad Eq. (1), coincides with $\Sigma^{\text{\tiny B}}_{t}(\hat{\sigma}_{z})$.
III (In)distinguishability of different unravelings of a master equation
We now turn to the question of whether the nonlinear quantities introduced earlier can be operationally used to distinguish between different unravelings. From the analysis in the previous section, as well as existing literature Wiseman and Diósi (2001); Caiaffa et al. (2017), it is evident that these quantities differentiate between unravelings of the same master equation. However, this distinction has so far been purely mathematical. The key question, therefore, is whether they also provide a means to experimentally distinguish between unravelings—i.e., whether they correspond to implementable procedures beyond their mathematical definition.
According to what described in the previous section and in Piñol et al. (2024), these nonlinear quantities are not measured but rather computed from the output signal, given knowledge of the stochastic master equation that models the measurement procedure. Specifically, they are reconstructed using the measurement record $y_{t}$ to update the conditional state $|\psi_{t}\rangle$, from which the conditional expectation $\langle\hat{O}\rangle_{t}$ and its powers can be computed for each realization of the noise. In this sense, they are not operational: the operational distinction arises at the level of implementing one measurement procedure over another—procedures that are different by definition. Consequently, the computed quantities merely quantify this predefined difference, arising from the choice of the measurement procedure.
To clarify this point, consider a scenario where a single von Neumann projective measurement replaces a continuous weak measurement. Suppose a spin-$\frac{1}{2}$ particle is initially in the maximally mixed state, $\hat{\rho}=I/2$, where $I$ is the identity matrix. If one performs a measurement of either $\hat{\sigma}_{z}$ or $\hat{\sigma}_{x}$, the measurement record in both cases is ±$\pm 1$. However, the conditional post-measurement state differs: in the first case, it is $|\uparrow_{z}\rangle$ or $|\downarrow_{z}\rangle$, while in the second, it is $|\uparrow_{x}\rangle$ or $|\downarrow_{x}\rangle$.
One can compute, for example, $\Sigma(\hat{\sigma}_{z})$, which amounts to 00 for a $\hat{\sigma}_{z}$ measurement and to $1$ for a $\hat{\sigma}_{x}$ measurement. While this quantity differs between the two cases and formally distinguishes them, the distinction was already known beforehand: the measurement procedures themselves, which are required to compute $\Sigma(\hat{\sigma}_{z})$, are different and naturally yield distinct results.
To be operational, nonlinear quantities should be measurable. Take, for example, $\langle\hat{O}\rangle_{t}$: by definition, this represents the expectation value of the observable $\hat{O}$ and, as such, can only be computed by repeating a measurement. However, these repeated measurements must be performed for a fixed realization of the noise, which is impossible.
The case is different for ${\mathbb{E}}_{\omega}[\langle\hat{O}^{2}\rangle_{\omega,t}]$ or $({\mathbb{E}}_{\omega}[\langle\hat{O}\rangle_{\omega,t}])^{2}$; the first quantity is mathematically equivalent to $\text{Tr}[\hat{A}^{2}\hat{\rho}_{t}]$, and the second to $(\text{Tr}[\hat{A}\hat{\rho}_{t}])^{2}$, where $\hat{\rho}_{t}$ is the averaged density matrix. In both cases, the quantum expectation is computed by performing a measurement of the statistical ensemble represented by $\hat{\rho}_{t}$, which is supposed to be accessible.
Another, more dramatic, consequence is that, if quantities like ${\mathbb{E}}_{\omega}[\langle\hat{O}\rangle^{2}_{\omega,t}]$ were accessible, one could establish a protocol for superluminal communication. To show this, let us consider a Bell-type experiment between two far-away observers, Alice and Bob, who share pairs of entangled $1/2$-spin particles in the singlet spin state:
[…]
(the spatial part of the wave function has been omitted). Suppose Alice decides to measure the spin of her particles along the $z$ direction; as a consequence, Bob will end up having a statistical mixtures of particles, half of which have spin up along $z$, and half spin down along the same direction. Then, effectively, Alice’s measurement, being random, amounts to a stochastic dynamics for Bob’s particles, which changes their initial state $|s\rangle$ into $|\psi\rangle_{\omega}$, where $\omega$ takes two values $\omega_{1}$ and $\omega_{2}$, each with probability $1/2$ and:
[…]
This is a simplified version of the dynamics discussed before in the context of the nonlinear unraveling given by Eq. (4) for a spin-$1/2$ particle. In this case, as already noted, $\Sigma(\hat{\sigma}_{z})=0$ and also ${\mathbb{E}}_{\omega}[\Sigma(\hat{\sigma}_{z})]=0$.
If on the other hand Alice decides to measure the spin of her particles along the $x$ direction, then Bob ends up having half of his particles with spin up and half with spin down along $x$. The random dynamics has the same structure as the previous one, with now:
[…]
This is a simplified version of the dynamical evolution given by the unraveling in Eq. (5), since the two components |$\ket{\uparrow_{z}}$ and |$\ket{\downarrow_{z}}$ receive different relative phases, depending on the realization of the noise. In this case $\Sigma(\hat{\sigma}_{z})=1$ and also ${\mathbb{E}}_{\omega}[\Sigma(\hat{\sigma}_{z})]=1$.
If ${\mathbb{E}}_{\omega}[\Sigma(\hat{\sigma}_{z})]$ were accessible to Bob, he and Alice would be able to establish a superluminal communication protocol. However, as discussed earlier, nonlinear quantities are not operationally accessible, and therefore, the superluminal protocol cannot be established.
The situation changes if Alice informs Bob of the measurement she performs; in this case, Bob can compute ${\mathbb{E}}_{\omega}[\Sigma(\hat{\sigma}_{z})]$. This does not violate relativity, as Alice and Bob must now establish a classical communication protocol. Moreover, Bob does not gain any additional knowledge beyond what is already contained in the measurement procedure disclosed by Alice.
IV Conclusion
There exist several stochastic unravelings for a given master equation. In light of this, a relevant question—both from a foundational and a practical perspective—is whether these unravelings can be distinguished. This issue can be addressed at both the mathematical and operational levels.
From a mathematical standpoint, it is well known that unravelings can be distinguished Wiseman and Diósi (2001); Caiaffa et al. (2017), not least because the very notion of different unravelings implies some form of differentiation. Thus, the real question is whether they can be distinguished operationally when they correspond to different measurement procedures.
As we have discussed, the answer is negative. Physical quantities are either unraveling-independent, such as $\mathbb{E}_{\omega}[\langle\hat{O}^{n}\rangle_{t}]$, and thus cannot be used to discriminate between unravelings; or they are unraveling-dependent, such as $\mathbb{E}_{\omega}[\langle\hat{O}\rangle^{n}_{t}]$ for $n>1$, but they are not operationally accessible unless the unraveling, or measurement procedure, which nonlinear quantities are supposed to distinguish, is already known.
V Acknowledgments
The authors wish to thank D. Keys and M. Paternostro for useful comments on the first draft of the paper. They acknowledge support from the EU EIC Pathfinder project QuCoM (101046973), INFN and the University of Trieste. A.B. acknowledges further support from the PNRR MUR projects PE0000023-NQSTI.
Appendix A Generalized Gaussian continuous measurements
We follow the approach of Ref. Jacobs (2014) to construct the measurement record and the corresponding Gaussian operators that lead to the family of unravelings in Eq. (2), so that we can interpret them as generalized Gaussian continuous quantum measurements of the observable $\hat{L}$.
Let us consider at a classical level a measurement of quantity $L$ that takes a continuum of values. As a consequence of the central limit theorem Fischer (2011), most measurements are expected to have a Gaussian error. Thus, the measurement result will be equal to the true value plus a Gaussian random variable. The extension to a quantum context, where one is measuring the observable $\hat{L}$, yields the following form of the measurement record
$\text{d}y_{t}=a\langle\hat{L}\rangle_{t}\text{d}t+b\frac{1}{2\sqrt{\gamma}}\text{d}W_{t},$ (21)
where $\langle\hat{L}\rangle_{t}=\bra{\psi_{t}}\hat{L}\ket{\psi_{t}}$ and $\text{d}W_{t}$ is a Wiener process. The constants $a$ and $b$ will be fixed in what follows. Now, let us consider the operator
$\hat{A}(\text{d}y_{t})=\mathcal{N}\exp\left(-\frac{\gamma}{\text{d}t}(c\text{d}y-d\hat{L}\text{d}t)^{2}\right),$ (22)
with $\mathcal{N}$ a suitable normalization constant and $c$ and $d$ a pair of constants to be determined. The structure of the operator in Eq. (9) can be thought as the quantum version of the likelihood function of a classical measurement of $L$ at time step $\text{d}t$. Applying this operator to the state |$\ket{\psi_{t}}$ and expanding up to order $\text{d}t$ yields
|$\ket{\psi_{t+\text{d}t}}\propto\left(1+2\gamma\alpha\langle\hat{L}\rangle\hat{L}-\gamma\beta\hat{L}^{2}\text{d}t+\sqrt{\gamma}\epsilon\hat{L}\text{d}W_{t}\right)\ket{\psi_{t}}$ (23)
where
$\alpha=cda,\qquad\beta=d^{2}\left(1-\frac{1}{2}c^{2}b^{2}\right),\qquad\epsilon=cdb,$ (24)
If we normalize this state, we obtain the following stochastic differential equation
$\begin{split}&\text{d}\ket{\psi_{t}}=\left(-\frac{\gamma}{2}(|\epsilon|^{2}-2\beta^{\text{\tiny{R}}})\langle\hat{L}^{2}\rangle_{t}\text{d}t-\gamma\beta\hat{L}^{2}\text{d}t\right.\\ &\left.+\gamma(2\alpha-\epsilon\epsilon^{\text{\tiny{R}}})\langle\hat{L}\rangle_{t}\hat{L}\text{d}t-\gamma\left(2\alpha^{\text{\tiny{R}}}-\frac{3}{2}(\epsilon^{\text{\tiny{R}}})^{2}\right)\langle\hat{L}\rangle_{t}^{2}\text{d}t\right.\\ &\left.+\sqrt{\gamma}(\epsilon\hat{L}-\epsilon^{\text{\tiny{R}}}\langle\hat{L}\rangle_{t})\text{d}W_{t}\right)\ket{\psi_{t}}.\end{split}$ (25)
The direct comparison of the above equation for |$\ket{\psi_{t}}$ with the family of unravelings in Eq. (2) leads to the following identifications
$\xi=\epsilon,\qquad|\xi|^{2}-2\beta^{\text{\tiny{R}}}=0,\qquad\beta^{\text{\tiny{I}}}=0,\qquad\alpha=\xi\xi^{\text{\tiny{R}}}.$ (26)
Solving for the parameters $a,b,c,d$, we find
$d^{\text{\tiny{R}}}=\sqrt{\frac{(\xi^{\text{\tiny{R}}})^{2}}{2}+\frac{1}{2}\sqrt{(\xi^{\text{\tiny{I}}}\xi^{\text{\tiny{R}}})^{2}+(\xi^{\text{\tiny{R}}})^{4}}},\qquad d^{\text{\tiny{I}}}=\frac{\xi^{\text{\tiny{R}}}\xi^{\text{\tiny{{I}}}}}{2d^{\text{\tiny{R}}}}.$ (27)
$a=b\xi^{\text{\tiny{R}}},\qquad c^{\text{\tiny{R}}}=\frac{\xi^{\text{\tiny{R}}}d^{\text{\tiny{R}}}+\xi^{\text{\tiny{I}}}d^{\text{\tiny{{I}}}}}{b|d|^{2}},\qquad c^{\text{\tiny{I}}}=\frac{\xi^{\text{\tiny{I}}}d^{\text{\tiny{R}}}-\xi^{\text{\tiny{R}}}d^{\text{\tiny{{I}}}}}{b|d|^{2}}$ (28)
To define the normalization constant $\mathcal{N}$, if we denote the measurement record by $\mathcal{A}=\text{d}y_{t}$ then we will require the expectation value of $\mathcal{A}$ to coincide with $\langle\hat{L}\rangle\text{d}t$. Thus, the relation
$\langle\mathcal{A}\rangle_{t}=\int_{-\infty}^{\infty}\mathcal{A}\text{Tr}[\hat{A}^{\dagger}(\mathcal{A})\hat{A}(\mathcal{A})]\text{d}\mathcal{A}=\langle\hat{X}\rangle_{t}\text{d}t$ (29)
must hold. Straightforward calculations show that this result can be met if the normalization constant $\mathcal{N}$ satisfies
$|\mathcal{N}|^{2}=\sqrt{\frac{2\gamma}{\pi\text{d}t}}\frac{((c^{\text{\tiny{R}}})^{2}-(c^{\text{\tiny{I}}})^{2})^{3/2}}{c^{\text{\tiny{R}}}d^{\text{\tiny{R}}}-c^{\text{\tiny{I}}}d^{\text{\tiny{I}}}}$ (30)
and if $\xi^{\text{\tiny{R}}}/b>0$. Therefore, it is enough to set $b=1$. From Eq. (21), we finally obtain the measurement record in Eq. (3).
A.1 Master equation and expectation values
Let us show that the family of unravelings in Eq. (2) lead to the density operator $\hat{\rho}_{t}$ that solves Eq. (1). Using Itô calculus, we have that the conditional state satisfies
$\begin{split}\text{d}\hat{\rho}_{\omega,t}&=(\text{d}\ket{\psi_{t}})\bra{\psi_{t}}+\ket{\psi_{t}}(\text{d}\bra{\psi_{t}})+(\text{d}\ket{\psi_{t}})(\text{d}\bra{\psi_{t}})\\ &=-\frac{i}{\hbar}[\hat{H},\hat{\rho}_{\omega,t}]\text{d}t-\frac{\lambda}{2}|\xi|^{2}[\hat{L},[\hat{L},\hat{\rho}_{\omega,t}]]\text{d}t\\ &+\sqrt{\lambda}(\xi^{\text{\tiny{R}}}(\{\hat{L},\hat{\rho}_{\omega,t}\}-2\langle\hat{L}\rangle_{t}\hat{\rho}_{\omega,t})+i\xi^{\text{\tiny{I}}}[\hat{L},\hat{\rho}_{\omega,t}])\text{d}W_{t}\end{split}$ (31)
Thus, averaging over all realizations, and using the constraint $|\xi|^{2}=1$, we see that $\hat{\rho}_{t}=\mathbb{E}_{\omega}[\hat{\rho}_{\omega,t}]$ satisfies
$\frac{\text{d}\hat{\rho}_{t}}{\text{d}t}=-\frac{i}{\hbar}[\hat{H},\hat{\rho}_{t}]-\frac{\lambda}{2}[\hat{L},[\hat{L},\hat{\rho}_{t}]],$ (32)
thus $\hat{\rho}_{t}$ is a solution of the master equation (1).
Now, let us consider a Hermitian operator $\hat{O}$. For the family of unravelings in Eq. (2), the evolution of the conditional expectation value of $\hat{O}$ reads
$\begin{split}&\text{d}\langle\hat{O}\rangle_{t}=-\frac{i}{\hbar}\langle[\hat{O},\hat{H}]\rangle_{t}\text{d}t-\frac{\lambda}{2}\langle[\hat{L},[\hat{L},\hat{O}]]\rangle_{t}\text{d}t\\ &+\sqrt{\lambda}(\xi^{\text{\tiny{R}}}(\langle\{\hat{O},\hat{L}\}\rangle_{t}-2\langle\hat{O}\rangle_{t}\langle\hat{L}\rangle_{t})+i\xi^{\text{\tiny{I}}}\langle[\hat{O},\hat{L}]\rangle_{t})\text{d}W_{t}\end{split}$ (33)
If follows immediately that the dynamical evolution of $\mathbb{E}_{\omega}[\langle\hat{O}\rangle_{t}]$ is the same for all the unravelings in Eq. (2), once the stochastic average is taken. This evolution coincides to that derived from the Lindblad Eq. (1), as required by construction.
Appendix B Free particle dynamics
Let us consider a free particle in a one-dimensional setting, and therefore fix the operators $\hat{H}$ and $\hat{L}$ as in Eq. (8).
Given the Gaussian ansatz in Eq. (9), one can write the quantum expectation values of observables of interest in terms of the Gaussian parameters as follows:
$\langle\hat{x}\rangle_{t}=\overline{x}_{t},\qquad\langle\hat{p}\rangle_{t}=\hbar\overline{k}_{t},\qquad\langle\hat{x}^{2}\rangle_{t}=\frac{1}{4a_{t}^{\text{\tiny{R}}}}+\overline{x}_{t}^{2},$ (34)
where $a_{t}^{\text{\tiny{R}}}$ is the real part of $a_{t}$. We will use these expressions to derive the dynamical evolution of $\langle\hat{x}\rangle^{2}_{t}$ for both unravelings and show that it is an unraveling-dependent quantity.
B.1 Nonlinear unraveling
Let us first consider the nonlinear unraveling of Eq. (6), and the Gaussian ansatz of Eq. (9). Using the results of Eq. (34), the dynamical evolution of ${\langle\hat{x}\rangle_{t}^{\text{\tiny{A}}}}^{2}$ is given by
$\text{d}{\langle\hat{x}\rangle_{t}^{\text{\tiny{A}}}}^{2}=\frac{2\hbar}{m}\overline{x}_{t}^{\text{\tiny{A}}}\overline{k}_{t}^{\text{\tiny{A}}}\text{d}t+\sqrt{\lambda}\overline{x}_{t}^{\text{\tiny{A}}}\frac{1}{a_{t}^{\text{\tiny{A}},\text{\tiny{R}}}}\text{d}W_{t}+\lambda\frac{1}{4{a_{t}^{\text{\tiny{A}},\text{\tiny{R}}}}^{2}}\text{d}t,$ (35)
Thus, we need to determine the evolution of the Gaussian parameters $\overline{x}_{t}^{\text{\tiny{A}}}$, $\overline{k}_{t}^{\text{\tiny{A}}}$ and $a_{t}^{{{\text{\tiny{A}}},\text{\tiny{R}}}}$, and we can do so by following the approach in Ref. Bassi (2005). With the choices for $\hat{H}$ and $\hat{O}$ of Eq. (8), the dynamical evolution in the position representation of the nonlinear unraveling in Eq. (4) reads
$\begin{split}{\text{d}{\psi_{t}^{\text{\tiny{A}}}}(x)}&=\left(\frac{i\hbar}{2m}\frac{\partial^{2}}{\partial x^{2}}\text{d}t-\frac{\lambda}{2}(x-\overline{x}_{t}^{\text{\tiny{A}}})^{2}\text{d}t\right.\\ &\left.+\sqrt{\lambda}(\hat{x}-\overline{x}_{t}^{\text{\tiny{A}}}){\text{d}W_{t}}\right){\psi_{t}^{\text{\tiny{A}}}}(x),\end{split}$ (36)
where we used Eq. (34). Recognizing that Gaussian states remain Gaussian, and after matching the terms on both sides of the above equation in powers of $x$, we obtain the following equations for the relevant parameters of the Gaussian state:
$\begin{split}{\text{d}a_{t}^{\text{\tiny{A}}}}&=\left(\lambda-\frac{2i\hbar}{m}{a_{t}^{\text{\tiny{A}}}}^{2}\right)\text{d}t,\\ {\text{d}\overline{x}_{t}^{\text{\tiny{A}}}}&=\frac{\hbar}{m}\overline{k}_{t}^{\text{\tiny{A}}}\text{d}t+\frac{\sqrt{\lambda}}{2a_{t}^{\text{\tiny{A}},\text{\tiny{R}}}}{\text{d}W_{t}},\\ {\text{d}\overline{k}_{t}^{\text{\tiny{A}}}}&=-\sqrt{\lambda}\frac{a_{t}^{\text{\tiny{A}},\text{\tiny{I}}}}{a_{t}^{\text{\tiny{A}},\text{\tiny{R}}}}{\text{d}W_{t}}.\end{split}$ (37)
Let us notice that the equation for $a_{t}^{\text{\tiny{A}}}$ is deterministic. The corresponding solution is reported in Ref. Bassi (2005). The real ($a_{t}^{\text{\tiny{A}},\text{\tiny{R}}}$) and imaginary ($a_{t}^{\text{\tiny{A}},\text{\tiny{I}}}$) parts of $a_{t}^{\text{\tiny{A}}}$ read
$\begin{split}a_{t}^{\text{\tiny{A}},\text{\tiny{R}}}&=\frac{c^{\text{\tiny{R}}}\sinh[2(b^{\text{\tiny{R}}}t+k^{\text{\tiny{R}}})]-c^{\text{\tiny{I}}}\sin[2(b^{\text{\tiny{I}}}t+k^{\text{\tiny{I}}})]}{\cosh[2(b^{\text{\tiny{R}}}t+k^{\text{\tiny{R}}})]+\cos[2(b^{\text{\tiny{I}}}t+k^{\text{\tiny{I}}})]},\\ a_{t}^{\text{\tiny{A}},\text{\tiny{I}}}&=\frac{c^{\text{\tiny{I}}}\sinh[2(b^{\text{\tiny{R}}}t+k^{\text{\tiny{R}}})+c^{\text{\tiny{R}}}\sin[2(b^{\text{\tiny{I}}}t+k^{\text{\tiny{I}}})]]}{\cosh[2(b^{\text{\tiny{R}}}t+k^{\text{\tiny{R}}})]+\cos[2(b^{\text{\tiny{I}}}t+k^{\text{\tiny{I}}})]},\end{split}$ (38)
where $b=(1+i)\sqrt{\hbar\lambda/m}$, $c=(1-i)/2\sqrt{m\lambda/\hbar}$, $k~{}=~{}\tanh^{-1}[a_{0}/c]$ and $a_{0}$ is the value of $a_{t}^{\text{\tiny{A}}}$ at time $t=0$.
Having solved for $a_{t}^{\text{\tiny{A}}}$, we can use the results of Eq. (37) to find the solution for $\overline{k}_{t}^{\text{\tiny{A}}}$. We obtain:
$\overline{k}_{t}^{\text{\tiny{A}}}=\overline{k}_{0}-\sqrt{\lambda}\int_{0}^{t}\frac{a_{s}^{\text{\tiny{A}},\text{\tiny{I}}}}{a_{s}^{\text{\tiny{A}},\text{\tiny{R}}}}\text{d}W_{s}.$ (39)
Finally, from the dynamical evolution for $\overline{x}_{t}^{\text{\tiny{A}}}$ in Eq. (37), the previous results for $a_{t}^{\text{\tiny{A}}}$ and $\overline{k}_{t}^{\text{\tiny{A}}}$ lead to:
$\begin{split}\overline{x}_{t}^{\text{\tiny{A}}}&=\overline{x}_{0}+\frac{\hbar}{m}\overline{k}_{0}t-\sqrt{\lambda}\frac{\hbar}{m}\int_{0}^{t}(t-s)\frac{a_{s}^{\text{\tiny{A}},\text{\tiny{I}}}}{a_{s}^{\text{\tiny{A}},\text{\tiny{R}}}}\text{d}W_{s}\\ &+\frac{\sqrt{\lambda}}{2}\int_{0}^{t}\frac{1}{a_{s}^{\text{\tiny{A}},\text{\tiny{R}}}}\text{d}W_{s}.\end{split}$ (40)
We are now in the position to solve Eq. (35):
$\begin{split}\mathbb{E}_{\omega}[({\langle\hat{x}\rangle_{t}^{\text{\tiny{A}}}})^{2}]&=\left(\overline{x}_{0}+\frac{\hbar}{m}\overline{k}_{0}t\right)^{2}\\ &+\lambda\int_{0}^{t}\text{d}s\frac{1}{(a_{s}^{\text{\tiny{A}},\text{\tiny{R}}})^{2}}\left(\frac{\hbar}{m}(t-s)a_{s}^{\text{\tiny{A}},\text{\tiny{I}}}-\frac{1}{2}\right)^{2},\end{split}$ (41)
which is the same result that one would obtain if by using the fact that $\langle\hat{x}\rangle_{t}^{\text{\tiny{A}}}=\overline{x}_{t}^{\text{\tiny{A}}}$ [cf. Eq. (34)], and square and average the result of Eq. (40).
As a final remark, from the fact that the variance $\text{Var}_{t}(\hat{x})$ is unraveling-independent, it follows from Eq. (12) and Eq. (10) that the term in the second line in Eq. (41) is equal to:
$\begin{split}&\lambda\int_{0}^{t}\text{d}s\frac{1}{(a_{s}^{\text{\tiny{A}},\text{\tiny{R}}})^{2}}\left(\frac{\hbar}{m}(t-s)a_{s}^{\text{\tiny{A}},\text{\tiny{I}}}-\frac{1}{2}\right)^{2}=\\ &=\frac{(m-2\hbar ta_{0}^{I})^{2}+4\hbar^{2}t^{2}(a_{0}^{R})^{2}}{4m^{2}a_{0}^{R}}+\lambda\frac{\hbar^{2}}{m^{2}}\frac{t^{3}}{3}+\\ &-\frac{1}{4}\frac{\cosh[2(b^{R}t+k^{R})]+\cos[2(b^{I}t+k^{I})]}{c^{I}\sinh[2(b^{R}t+k^{R})+c^{R}\sin[2(b^{I}t+k^{I})]]}.\end{split}$ (42)
B.2 Linear unraveling
Let us now consider the linear unraveling of Eq. (5), with the same Gaussian ansatz of Eq. (9). In this case, the results of Eq. (34) lead to the following evolution for ${\langle\hat{x}\rangle_{t}^{\text{\tiny{B}}}}^{2}$:
$\text{d}{\langle\hat{x}\rangle_{t}^{\text{\tiny{B}}}}^{2}=\frac{2\hbar}{m}\overline{x}_{t}^{\text{\tiny{B}}}\overline{k}_{t}^{\text{\tiny{B}}}\text{d}t.$ (43)
Working in the position representation, the evolution of the wave function corresponds to:
${\text{d}{\psi_{t}^{\text{\tiny{B}}}}(x)}=\left(\frac{i\hbar}{2m}\frac{\partial^{2}}{\partial x^{2}}{\text{d}t}-\frac{\lambda}{2}x^{2}\text{d}t-i\sqrt{\lambda}{x}{\text{d}W_{t}}\right){\psi_{t}^{\text{\tiny{B}}}(x)},$ (44)
and the dynamical evolution for the relevant parameters of the Gaussian state reads:
$\begin{split}{\text{d}a_{t}^{\text{\tiny{B}}}}&=-\frac{2i\hbar}{m}{a_{t}^{\text{\tiny{B}}}}^{2}{\text{d}t},\\ {\text{d}\overline{x}_{t}^{\text{\tiny{B}}}}&=\frac{\hbar}{m}\overline{k}_{t}^{\text{\tiny{B}}}{\text{d}t},\\ {\text{d}\overline{k}_{t}^{\text{\tiny{B}}}}&=-\sqrt{\lambda}{\text{d}W_{t}}.\end{split}$ (45)
Solving for $\overline{k}_{t}^{\text{\tiny{B}}}$, we obtain:
$\overline{k}_{t}^{\text{\tiny{B}}}=\overline{k}_{0}-\sqrt{\lambda}\int_{0}^{t}\text{d}W_{s},$ (46)
and the substitution of this result in the evolution of $\overline{x}_{t}^{\text{\tiny{B}}}$ in Eq. (45) yields to:
$\overline{x}_{t}^{\text{\tiny{B}}}=\overline{x}_{0}+\frac{\hbar}{m}\overline{k}_{0}t-\frac{\hbar\sqrt{\lambda}}{m}\int_{0}^{t}(t-s)\text{d}W_{s},$ (47)
with solution:
$\begin{split}\mathbb{E}_{\omega}[{\langle\hat{x}\rangle_{t}^{\text{\tiny{B}}}}^{2}]=\left(\overline{x}_{0}+\frac{\hbar}{m}\overline{k}_{0}t\right)^{2}+\lambda\frac{\hbar^{2}}{m^{2}}\int_{0}^{t}(t-s)^{2}\text{d}s,\end{split}$ (48)
which in general differs from the expression for $\mathbb{E}_{\omega}[{\langle\hat{x}\rangle_{t}^{\text{\tiny{A}}}}^{2}]$, as shown in Fig. 3. Finally, solving for $a_{t}^{\text{\tiny{B}}}$, we obtain
$\begin{split}a_{t}^{\text{\tiny{B}},\text{\tiny{R}}}&=\frac{m^{2}a_{0}^{R}}{(m-2\hbar ta_{0}^{I})^{2}+4\hbar^{2}t^{2}(a_{0}^{R})^{2}},\\ a_{t}^{\text{\tiny{B}},\text{\tiny{I}}}&=\frac{m^{2}a_{0}^{I}-2\hbar t|a_{0}|^{2}}{(m-2\hbar ta_{0}^{I})^{2}+4\hbar^{2}t^{2}(a_{0}^{R})^{2}}.\end{split}$ (49)
Figure 3: Stochastic average over all realizations of ${\langle\hat{x}\rangle_{t}}^{2}$, for the nonlinear (${\langle\hat{x}\rangle_{t}^{\text{\tiny{A}}}}^{2}$ in Eq. (41)) and linear (${\langle\hat{x}\rangle_{t}^{\text{\tiny{B}}}}^{2}$ in Eq. (48)) unravelings. For illustrative purposes, we use the parameters: $\overline{x}_{0}~{}=~{}0$, $\overline{k}_{0}~{}=~{}0$, $m~{}=~{}10^{-15}\mathrm{kg}$, $a_{0}~{}=~{}0.25\times 10^{9}\mathrm{m^{2}}$, $\lambda~{}=~{}10^{23}\mathrm{m^{-2}\,Hz}$
Appendix C Particle in a harmonic trap
Let us now consider a particle trapped in a harmonic potential:
$\hat{H}=\frac{1}{2m}\hat{p}^{2}+\frac{1}{2}m\Omega^{2}\hat{x}^{2},$ (50)
and we maintain the choice $\hat{L}=\hat{x}$. We derive the dynamical evolution for both unravelings under the Gaussian ansatz of Eq. (9).
C.1 Nonlinear unraveling
We can follow the same procedure as in the previous section; the parameters of the Gaussian state $a_{t}^{\text{\tiny{A}}}$, $\overline{x}_{t}^{\text{\tiny{A}}}$ and $\overline{k}_{t}^{\text{\tiny{A}}}$ now obey the following dynamical evolution:
$\begin{split}\text{d}a_{t}^{\text{\tiny{A}}}&=\left(\lambda+i\frac{m\Omega^{2}}{2\hbar}-\frac{2i\hbar}{m}{a_{t}^{\text{\tiny{A}}}}^{2}\right)\text{d}t,\\ {\text{d}\overline{x}_{t}^{\text{\tiny{A}}}}&=\frac{\hbar}{m}\overline{k}_{t}^{\text{\tiny{A}}}\text{d}t+\frac{\sqrt{\lambda}}{2a_{t}^{\text{\tiny{A}},\text{\tiny{R}}}}{\text{d}W_{t}},\\ {\text{d}\overline{k}_{t}^{\text{\tiny{A}}}}&=-\frac{m\Omega^{2}}{\hbar}\overline{x}_{t}^{\text{\tiny{A}}}\text{d}t-\sqrt{\lambda}\frac{a_{t}^{\text{\tiny{A}},\text{\tiny{I}}}}{a_{t}^{\text{\tiny{A}},\text{\tiny{R}}}}{\text{d}W_{t}}.\end{split}$ (51)
The solutions for $\overline{x}_{t}^{\text{\tiny{A}}}$ and $\overline{k}_{t}^{\text{\tiny{A}}}$ read:
$\begin{split}&\overline{x}_{t}^{\text{\tiny{A}}}=\frac{\hbar}{m\Omega}\overline{k}_{0}\sin\left(\Omega t\right)+\overline{x}_{0}\cos\left(\Omega t\right)+\\ &+\sqrt{\lambda}\int_{0}^{t}\frac{2a_{t}^{\text{\tiny{A}},\text{\tiny{I}}}\hbar\sin\left(\Omega(s-t)\right)+m\Omega\cos\left(\Omega(s-t)\right)}{2a_{t}^{\text{\tiny{A}},\text{\tiny{R}}}m\Omega}\,\text{d}W_{s},\\ &\overline{k}_{t}^{\text{\tiny{A}}}=-\frac{m\Omega}{\hbar}\overline{x}_{0}\sin(\Omega t)+\overline{k}_{0}\cos\left(\Omega t\right)+\\ &+\sqrt{\lambda}\int_{0}^{t}\frac{m\Omega\sin\left(\Omega(s-t)\right)-2a_{t}^{\text{\tiny{A}},\text{\tiny{I}}}\hbar\cos\left(\Omega(s-t)\right)}{2a_{t}^{\text{\tiny{A}},\text{\tiny{R}}}\hbar}\,\text{d}W_{s}.\end{split}$ (52)
Thus, taking the square of $\overline{x}_{t}$ one gets:
[…]
The solution for $a_{t}^{\text{\tiny{A}}}$ is as in Eq. (38), where the parameters $b$ and $c$ are now defined as:
$b=\frac{\hbar}{m}\sqrt{2(\sqrt{x_{a}^{2}+y_{a}^{2}}-x_{a})}+i\frac{\hbar}{m}\sqrt{2\left(\sqrt{x_{a}^{2}+y_{a}^{2}}+x_{a}\right)},$ (54)
with $x_{a}=m^{2}\Omega^{2}/(4\hbar^{2})$ and $y_{a}=m\lambda/(2\hbar)$. On the other hand, we have
$c=\sqrt{\frac{1}{2}\left(\sqrt{x_{a}^{2}+y_{a}^{2}}+x_{a}\right)}-i\sqrt{\frac{1}{2}\left(\sqrt{x_{a}^{2}+y_{a}^{2}}-x_{a}\right)}.$ (55)
We notice that in the limit in which $\Omega\rightarrow 0$, we have that $x_{a}\rightarrow 0$ and thus we recover the definitions of the parameters $c$ and $b$ given after Eq. (38).
C.2 Linear unraveling
The stochastic differential equations for the Gaussian state parameters $a_{t}^{\text{\tiny{B}}}$, $\overline{x}_{t}^{\text{\tiny{B}}}$ and $\overline{k}_{t}^{\text{\tiny{B}}}$ are given by:
$\begin{split}{\text{d}a_{t}^{\text{\tiny{B}}}}&=\left(i\frac{m\Omega^{2}}{2\hbar}-\frac{2i\hbar}{m}{a_{t}^{\text{\tiny{B}}}}^{2}\right)\text{d}t,\\ {\text{d}\overline{x}_{t}^{\text{\tiny{B}}}}&=\frac{\hbar}{m}\overline{k}_{t}^{\text{\tiny{B}}}{\text{d}t},\\ {\text{d}\overline{k}_{t}^{\text{\tiny{B}}}}&=-\frac{m\Omega^{2}}{\hbar}\overline{x}_{t}^{\text{\tiny{B}}}{\text{d}t}-\sqrt{\lambda}{\text{d}W_{t}}.\end{split}$ (56)
The above equations lead to the following solutions for $\overline{x}_{t}^{\text{\tiny{B}}}$ and $\overline{k}_{t}^{\text{\tiny{B}}}$:
$\begin{split}&\overline{x}_{t}^{\text{\tiny{B}}}\!\!=\!\!\frac{\hbar}{m\Omega}\overline{k}_{0}\sin(\Omega t)\!+\!\overline{x}_{0}\cos\left(\Omega t\right)\!+\!\frac{\sqrt{\lambda}\hbar}{m\Omega}\!\!\int_{0}^{t}\!\!\sin(\Omega(s-t))\text{d}W_{s},\\ &\overline{k}_{t}^{\text{\tiny{B}}}\!\!=\!\!-\frac{m\Omega}{\hbar}\overline{x}_{0}\sin(\Omega t)\!+\!\overline{k}_{0}\cos\left(\Omega t\right)\!-\!\sqrt{\lambda}\!\int_{0}^{t}\!\!\cos\left(\Omega(s-t)\right)\!\text{d}W_{s},\end{split}$ (57)
and it follows that:
$\begin{split}&\mathbb{E}_{\omega}[{\langle\hat{x}\rangle_{t}^{\text{\tiny{B}}}}^{2}]=\left(\frac{\hbar}{m\Omega}\overline{k}_{0}\sin\left(\Omega t\right)+\overline{x}_{0}\cos\left(\Omega t\right)\right)^{2}+\\ &+\int_{0}^{t}\left(\frac{\sqrt{\lambda}\hbar\sin\left(\Omega(s-t)\right)}{m\Omega}\right)^{2}\,\text{d}s.\end{split}$ (58)
We notice that the equation for $a_{t}^{\text{\tiny{B}}}$ is the particular case $\lambda=0$ of Eq. (51). Therefore, the solution of this equation can be readily obtained from Eq. (38), where we set $y_{a}=0$ and find
$b=i\Omega,\qquad c=\frac{m\Omega}{2\hbar}.$ (59)
Figure 4: Time evolution of the unraveling-independent variance $\text{Var}_{t}(\hat{x})$ (Eq. (64), green line), and of $\Sigma_{t}(\hat{x})$ for the nonlinear unraveling ($\Sigma_{t}^{\text{\tiny{A}}}(\hat{x})$ as in Eq. (10), with $b$ and $c$ given in Eq. (54) and Eq. (55) [blue line]) and the linear unraveling ($\Sigma_{t}^{\text{\tiny{B}}}(\hat{x})$ as in Eq. (10), with $b$ and $c$ given in Eq. (59) [yellow line]). For illustrative purposes, we use the parameters: $\overline{x}_{0}~{}=~{}0$, $\overline{k}_{0}~{}=~{}0$, $m~{}=~{}10^{-15}\,\mathrm{kg}$, $a_{0}~{}=~{}10^{-3}\,\mathrm{m^{2}}$, $\lambda~{}=~{}10^{23}\,\mathrm{m^{-2}\,Hz}$. The three quantities evolve differently over time. In addition here we assume $\Omega=10^{4}\,\mathrm{Hz}$ for the harmonic potential term.
C.3 Covariance matrix and the Quantum Riccati equation
As mentioned in the Introduction, for optomechanical systems it is costumary to work with the components of the covariance matrix $\Sigma_{t}$, given by
$\Sigma_{t}=\left(\begin{array}[]{cc}\Sigma_{t}(\hat{x})&\Sigma_{t}(\hat{x},\hat{p})\\ \Sigma_{t}(\hat{x},\hat{p})&\Sigma_{t}(\hat{p})\end{array}\right).$ (60)
We again restrict the analysis to Gaussian states. In this case, $\Sigma_{t}$ satisfies a quantum Riccati equation Genoni et al. (2016):
$\frac{\text{d}\Sigma_{t}}{\text{d}t}=\alpha\Sigma_{t}+\Sigma_{t}\alpha^{T}+\delta-\Sigma_{t}\beta\beta^{T}\Sigma_{t},$ (61)
where $\alpha,\beta$ and $\delta$ are the drift, diffusion and backaction matrices, respectively Wu et al. (2024). We are interested in the first component $\Sigma_{t}(\hat{x})$.
For the nonlinear unraveling, $\Sigma_{t}^{\text{\tiny{A}}}(\hat{x})$ is as in Eq. (10), with the constants $b$ and $c$ as defined in Eq. (54) and Eq. (55), respectively. As a matter of fact, the covariance matrix $\Sigma_{t}^{\text{\tiny{A}}}$ satisfies Eq. (61), with
$\alpha^{\text{\tiny{A}}}\!=\!\left(\begin{array}[]{cc}0&\frac{1}{m}\\ -m\Omega^{2}&0\end{array}\right)\!,\,\,\beta^{\text{\tiny{A}}}\!=\!\left(\begin{array}[]{cc}0&2\sqrt{\lambda}\\ 0&0\end{array}\right)\!,\,\,\delta^{\text{\tiny{A}}}\!=\!\left(\begin{array}[]{cc}0&0\\ 0&\lambda\hbar^{2}\end{array}\right)\!.$ (62)
This is the Riccati equation typically used in the literature Genoni et al. (2016); Wu et al. (2024); Piñol et al. (2024); Albarelli et al. (2017); Magrini et al. (2021); Zhang and Mølmer (2017); Belenchia et al. (2020); Zhang et al. (2017), which often contains additional contributions coming from other noise sources.
In the case of the linear unraveling, $\Sigma_{t}^{\text{\tiny{B}}}(\hat{x})$ is as $\Sigma_{t}^{\text{\tiny{A}}}(\hat{x})$, but with the constants $b$ and $c$ given in Eq. (59). As one might have expected, the Riccati equation satisfied by $\Sigma_{t}^{\text{\tiny{B}}}$ differs from the one for $\Sigma_{t}^{\text{\tiny{A}}}$, as in this case:
$\alpha^{\text{\tiny{B}}}\!=\!\left(\begin{array}[]{cc}0&\frac{1}{m}\\ -m\Omega^{2}&0\end{array}\right),\qquad\beta^{\text{\tiny{B}}}=0_{2\times 2},\qquad\delta^{\text{\tiny{B}}}=0_{2\times 2}.$ (63)
Finally, the unraveling-independent variance $\text{Var}_{t}(\hat{x})$ is given by
$\begin{split}\text{Var}_{t}(\hat{x})&=\frac{\hbar}{2m\Omega}\frac{\cosh[2k^{\text{\tiny{R}}}]+\cos[2(\Omega t+k^{\text{\tiny{I}}})]}{\sinh[2k^{\text{\tiny{R}}}]}\\ &+\frac{\lambda\hbar^{2}}{2m^{2}\Omega^{2}}\left(t-\frac{\sin(2\Omega t)}{2\Omega}\right),\end{split}$ (64)
with
$k^{\text{\tiny{R}}}=\frac{1}{4}\ln(x_{v}^{2}+y_{v}^{2}),\qquad k^{\text{\tiny{I}}}=\frac{1}{2}\arctan\left(\frac{y_{v}}{x_{v}}\right),$ (65)
where
$\begin{split}x_{v}&=\frac{|c|^{2}-|a_{0}|^{2}}{(c^{\text{\tiny{R}}}-a_{0}^{\text{\tiny{R}}})^{2}+(c^{\text{\tiny{I}}}-a_{0}^{\text{\tiny{I}}})^{2}},\\ y_{v}&=\frac{2(a_{0}^{\text{\tiny{I}}}c^{\text{\tiny{R}}}-a_{0}^{\text{\tiny{R}}}c^{\text{\tiny{I}}})}{(c^{\text{\tiny{R}}}-a_{0}^{\text{\tiny{R}}})^{2}+(c^{\text{\tiny{I}}}-a_{0}^{\text{\tiny{I}}})^{2}}.\end{split}$ (66)
The covariance matrix $\text{Var}_{t}$, defined in analogy with Eq. (60) also satisfies a quantum Riccati equation (61), where:
$\alpha^{\text{\tiny{Var}}}\!=\!\left(\begin{array}[]{cc}0&\frac{1}{m}\\ -m\Omega^{2}&0\end{array}\right)\!,\,\,\beta^{\text{\tiny{Var}}}=0_{2\times 2},\,\,\delta^{\text{\tiny{Var}}}\!=\!\left(\begin{array}[]{cc}0&0\\ 0&\lambda\hbar^{2}\end{array}\right)\!,$ (67)
which differs from those of the two unravelings considered. We remark that the Riccati equation for $\text{Var}_{t}$ is unraveling-independent.
Appendix D Spin-$1/2$ particle - solution of the nonlinear unraveling
In order to find the solution for the wavefunction |$\ket{\psi_{t}^{\text{\tiny{A}}}}$ of the nonlinear unraveling of Eq. (4) for our spin-$1/2$ system, let us consider the following linear unraveling:
$\text{d}\ket{\phi_{t}^{\text{\tiny{A}}}}=\left(-i\nu\hat{\sigma}_{z}\text{d}t-\frac{\lambda}{2}\hat{\sigma}_{z}^{2}\text{d}t+\sqrt{\lambda}\hat{\sigma}_{z}\text{d}\xi_{t}\right)\ket{\phi_{t}^{\text{\tiny{A}}}}$ (68)
where $\xi_{t}$ is a Wiener process. A straightforward application of Itô calculus leads to the following stochastic differential equation (SDE) for $\lVert\phi_{t}^{\text{\tiny{A}}}\rVert^{2}=\bra{\phi_{t}^{\text{\tiny{A}}}}\phi_{t}^{\text{\tiny{A}}}\rangle$:
${\text{d}}\lVert\phi_{t}^{\text{\tiny{A}}}\rVert^{2}=2\sqrt{\lambda}\bra{\phi_{t}^{\text{\tiny{A}}}}\hat{\sigma}_{z}\ket{\phi_{t}^{\text{\tiny{A}}}}\text{d}\xi_{t}$ (69)
and from this it follows that:
$\text{d}\lVert\phi_{t}^{\text{\tiny{A}}}\rVert^{-1}=-\frac{1}{\lVert\phi_{t}^{\text{\tiny{A}}}\rVert}\sqrt{\lambda}\langle\hat{\sigma}_{z}\rangle_{t}^{\text{\tiny{A}}}\text{d}\xi_{t}+\frac{3}{2}\frac{1}{\lVert\phi_{t}^{\text{\tiny{A}}}\rVert}\lambda{\langle\hat{\sigma}_{z}\rangle_{t}^{\text{\tiny{A}}}}^{2}\text{d}t,$ (70)
where $\langle\hat{\sigma}_{z}\rangle_{t}^{\text{\tiny{A}}}=\bra{\phi_{t}^{\text{\tiny{A}}}}\hat{\sigma}_{z}\ket{\phi_{t}^{\text{\tiny{A}}}}/\lVert\phi_{t}^{\text{\tiny{A}}}\rVert$. From these equations, it follows that if we define |$\ket{\psi_{t}^{\text{\tiny{A}}}}=\ket{\phi_{t}^{\text{\tiny{A}}}}/\lVert\phi_{t}^{\text{\tiny{A}}}\rVert$, we have that |$\ket{\psi_{t}^{\text{\tiny{A}}}}$ satisfies the following SDE:
$\begin{split}\text{d}\ket{\psi_{t}^{\text{\tiny{A}}}}=&\left(-i\nu\hat{\sigma}_{z}\text{d}t+\sqrt{\lambda}(\hat{\sigma}_{z}-\langle\hat{\sigma}_{z}\rangle_{t}^{\text{\tiny{A}}})\right.\\ &\left.-\frac{\lambda}{2}(\hat{\sigma}_{z}^{2}+2\hat{\sigma}_{z}\langle\hat{\sigma}_{z}\rangle_{t}^{\text{\tiny{A}}}-3{\langle\hat{\sigma}_{z}\rangle_{t}^{\text{\tiny{A}}}}^{2})\text{d}t\right)\ket{\psi_{t}^{\text{\tiny{A}}}}\end{split}$ (71)
Let us consider now the following change of measure Gatarek and Gisin (1991); Bassi (2005):
$\text{d}\xi_{t}=\text{d}W_{t}+2\sqrt{\lambda}\langle\hat{\sigma}_{z}\rangle_{t}^{\text{\tiny{A}}}\text{d}t,$ (72)
we have:
$\text{d}\ket{\psi_{t}^{\text{\tiny{A}}}}\!\!=\!\!\left(-{i}\nu\hat{\sigma}_{z}\text{d}t+\sqrt{\lambda}(\hat{\sigma}_{z})_{t}^{\text{\tiny{A}}}\text{d}W_{t}-\frac{\lambda}{2}{(\hat{\sigma}_{z})_{t}^{\text{\tiny{A}}}}^{2}\text{d}t\right)\!\!\ket{\psi_{t}^{\text{\tiny{A}}}},$ (73)
which is Eq. (4) for our spin-1/2 model.
Therefore, the above calculations show that in order to solve for |$\ket{\psi_{t}^{\text{\tiny{A}}}}$, it suffices to solve for |$\ket{\phi_{t}^{\text{\tiny{A}}}}$, normalize, and perform the change of measure of Eq. (72). Therefore, solving for |$\ket{\phi_{t}^{\text{\tiny{A}}}}$, one finds:
|$\ket{\phi_{t}^{\text{\tiny{A}}}}=\exp\left[-{i}\nu\hat{\sigma}_{z}t-\lambda\hat{\sigma}_{z}^{2}t+\sqrt{\lambda}\hat{\sigma}_{z}\xi(t)\right]\ket{\psi_{0}},$ (74)
with |$\ket{\psi_{0}}$ the initial state of the system, which we assume to be normalized. After normalizing and performing the change of measure of Eq. (72), we obtain Eq. (14).
Appendix E Collapse in the nonlinear unraveling
Let us consider the nonlinear unraveling in Eq. (4) for the case in which $\hat{O}=\hat{L}=(\hbar\nu)^{-1}\hat{H}=\hat{\sigma}_{z}$. In this case, one can directly make the identification $\sqrt{\lambda}=\frac{1}{2}\sigma$ in Eq. (11) of the work of S. Adler et al. in Ref. Adler et al. (2001), and calculate the dynamical evolution for $\langle\hat{\sigma}_{z}^{n}\rangle_{t}^{\text{\tiny{A}}}$ ($n\in\mathbb{N})$, yielding
$\text{d}\langle\hat{\sigma}_{z}^{n}\rangle_{t}^{\text{\tiny{A}}}=2\sqrt{\lambda}(\langle\hat{\sigma}_{z}^{n+1}\rangle_{t}^{\text{\tiny{A}}}-\langle\hat{\sigma}_{z}\rangle_{t}^{\text{\tiny{A}}}\langle\hat{\sigma}_{z}^{n}\rangle_{t}^{\text{\tiny{A}}})\text{d}W_{t},$ (75)
and in particular for the variance process $\Sigma^{\text{\tiny{A}}}_{t}(\hat{\sigma}_{z})~{}=~{}\langle(\hat{\sigma}_{z}~{}-~{}\langle\hat{\sigma}_{z}\rangle_{t}^{\text{\tiny{A}}})^{2}\rangle_{t}^{\text{\tiny{A}}}$, one obtains the following evolution
$\text{d}\Sigma^{\text{\tiny{A}}}_{t}(\hat{\sigma}_{z})=-4\lambda{\Sigma_{t}^{\text{\tiny{A}}}}^{2}(\hat{\sigma}_{z})\text{d}t+2\sqrt{\lambda}\beta_{t}\text{d}W_{t},$ (76)
with $\beta_{t}$ is the third central moment of $\hat{\sigma}_{z}$. With these considerations, in Ref. Adler et al. (2001) it is shown that $\Sigma_{t}^{A}(\hat{\sigma}_{z})$ is a supermartingale. Moreover, one arrives at
$\mathbb{E}_{\omega}[\Sigma^{\text{\tiny{A}}}_{t}(\hat{\sigma}_{z})]\leq\frac{\Sigma^{\text{\tiny{A}}}_{0}(\hat{\sigma}_{z})}{1+4\lambda\Sigma^{\text{\tiny{A}}}_{0}(\hat{\sigma}_{z})t},$ (77)
which implies that in the limit $t\rightarrow\infty$, $\mathbb{E}_{\omega}[\Sigma_{t}^{A}]$ vanishes. Since the process itself is non-negative, then Eq. (15) follows almost surely.

### v2

@@
 (1993); Daley (2014); Diósi (1998); Ghirardi et al.
-(1990); Kafri et al.
+(1990a); Kafri et al.
 (2014); Tilloy and Diósi (2016, 2017); Weiss (2012); Hänggi et al.
-(1990); Wiseman and Diósi (2001); Caiaffa et al.
+(1990); Ghirardi et al.
+(1990b); Pearle (1997); Wiseman and Diósi (2001); Adler and Bassi (2007); Caiaffa et al.
 (2017); Jacobs (2014); Wiseman (1996); Carmichael (2007); Piccitto et al.
@@
 Stochastic Schrödinger equations are also used in the foundations of quantum mechanics to model the spontaneous collapse of the wave function Diósi (1998); Ghirardi et al.
-(1990); more recently, they have been use to model the coupling between classical gravity and quantum matter Kafri et al.
+(1990a); more recently, they have been use to model the coupling between classical gravity and quantum matter Kafri et al.
 (2014); Tilloy and Diósi (2016, 2017) offering an alternative to quantum gravity.
-The association between master equations and stochastic unravelings is not unique: in general, there are infinite ways to unravel a master equation via SSEs Wiseman and Diósi (2001); Caiaffa et al.
+The association between master equations and stochastic unravelings is not unique: in general, there are infinite ways to unravel a master equation via SSEs Ghirardi et al.
+(1990b); Pearle (1997); Wiseman and Diósi (2001); Adler and Bassi (2007); Caiaffa et al.
 (2017), as also shown later on.
@@
 This question is both interesting and relevant to the theory of open quantum systems.
-However, here we show that this is not the case.
-The reason is that nonlinear quantities are either reconstructed from the unraveling associated with the measurement procedure, and thus quantify differences that are embodied in the choice of the unraveling—i.e., the operational distinction is already made in selecting the measurement procedure, not in computing the associated nonlinear quantities—or they are not accessible at all.
+We argue that nonlinear quantities are either reconstructed from the unraveling associated with the measurement procedure, and thus quantify differences that are embodied in the choice of the unraveling—i.e., the operational distinction is already made in selecting the measurement procedure, not in computing the associated nonlinear quantities—or they are not accessible at all.
 The latter point is particularly crucial, as their accessibility would enable the development of a protocol for superluminal signaling, which is (supposed to be) forbidden.
@@
 Throughout this work we will use the Itô formalism.
-With the exception of the limiting case $\xi^{\text{\tiny R}}=0$, one can show  [^2] See Appendix A for the details that the above family of unravelings corresponds to different types of Gaussian continuous measurement Jacobs (2014), with measurement record
+With the exception of the limiting case $\xi^{\text{\tiny R}}=0$, one can show  [^2] See Appendix A for the details.
+that the above family of unravelings corresponds to different types of Gaussian continuous measurement Jacobs (2014), with measurement record
 $\text{d}y_{t}=\xi^{\text{\tiny{R}}}\langle\hat{L}\rangle_{t}\text{d}t+\frac{1}{2\sqrt{\gamma}}\text{d}W_{t}.$ (3)
@@
 is solution for both unravelings, where $a_{t}$ and $\gamma_{t}$ are complex functions of time, whereas $\overline{x}_{t}$ and $\overline{k}_{t}$ are real.
-It can be shown  [^4] See Appendix B, where we present the explicit time dependence of $a_{t}$, $\gamma_{t}$, $\overline{x}_{t}$, $\overline{k}_{t}$, and we derive the free particle dynamics in the different unravelings that for the nonlinear unraveling of Eq.
+It can be shown  [^4] See Appendix B, where we present the explicit time dependence of $a_{t}$, $\gamma_{t}$, $\overline{x}_{t}$, $\overline{k}_{t}$, and we derive the free particle dynamics in the different unravelings.
+that for the nonlinear unraveling of Eq.
 (4), with the choices in Eq.
@@
 $\Sigma^{\text{\tiny A}}_{t}(\hat{x})=\frac{1}{4}\frac{\cosh[2(b^{\text{\tiny{R}}}t+k^{\text{\tiny{R}}})]+\cos[2(b^{\text{\tiny{I}}}t+k^{\text{\tiny{I}}})]}{c^{\text{\tiny{R}}}\sinh[2(b^{\text{\tiny{R}}}t+k^{\text{\tiny{R}}})]-c^{\text{\tiny{I}}}\sin[2(b^{\text{\tiny{I}}}t+k^{\text{\tiny{I}}})]},$ (10)
-with $b,c$ and $k$ constants 555cf. Eq. (38). On the other hand, for the linear unraveling, we obtain:
+with $b,c$ and $k$ constants 555cf. Eq. (38).. On the other hand, for the linear unraveling, we obtain:
 $\Sigma^{\text{\tiny B}}_{t}(\hat{x})=\frac{1}{4}\frac{(m-2\hbar ta_{0}^{\text{\tiny{I}}})^{2}+4\hbar^{2}t^{2}(a_{0}^{\text{\tiny{R}}})^{2}}{m^{2}a_{0}^{\text{\tiny{R}}}}.$ (11)
 Note that in both cases $\Sigma_{t}(\hat{x})$ does not depend on the noise, therefore the stochastic average is ineffective.
-Both quantities differ from the variance $\text{Var}_{t}(\hat{x})=\mathbb{E}_{\omega}[\langle\hat{x}^{2}\rangle_{t}]-(\mathbb{E}_{\omega}[\langle\hat{x}\rangle])^{2}_{t}=\text{Tr}[\hat{x}^{2}\hat{\rho}_{t}]-(\text{Tr}[\hat{x}\hat{\rho}_{t}])^{2}$  [^6] In general, we have $\mathbb{E}_{\omega}[\text{Var}(\hat{O}_{i},\hat{O}_{j})]=\frac{1}{2}\mathbb{E}_{\omega}\langle\{\hat{O}_{i},\hat{O}_{j}\}\rangle_{t}-\mathbb{E}_{\omega}[\langle\hat{O}_{i}\rangle_{t}]\mathbb{E}_{\omega}[\langle\hat{O}_{j}\rangle_{t}]$, obtained from the density matrix, which by construction is unraveling-independent.
+Both quantities differ from the variance $\text{Var}_{t}(\hat{x})=\mathbb{E}_{\omega}[\langle\hat{x}^{2}\rangle_{t}]-(\mathbb{E}_{\omega}[\langle\hat{x}\rangle])^{2}_{t}=\text{Tr}[\hat{x}^{2}\hat{\rho}_{t}]-(\text{Tr}[\hat{x}\hat{\rho}_{t}])^{2}$  [^6] In general, we have $\mathbb{E}_{\omega}[\text{Var}(\hat{O}_{i},\hat{O}_{j})]=\frac{1}{2}\mathbb{E}_{\omega}\langle\{\hat{O}_{i},\hat{O}_{j}\}\rangle_{t}-\mathbb{E}_{\omega}[\langle\hat{O}_{i}\rangle_{t}]\mathbb{E}_{\omega}[\langle\hat{O}_{j}\rangle_{t}]$., obtained from the density matrix, which by construction is unraveling-independent.
 Straightforward calculations show that
@@
 For the nonlinear unraveling given by Eq.
-(4), the solution is  [^7] See Appendix D, where we derive an explicit solution for the Spin-$1/2$ particle in the nonlinear unraveling:
+(4), the solution is  [^7] See Appendix D, where we derive an explicit solution for the Spin-$1/2$ particle in the nonlinear unraveling.:
 |$\ket{\psi^{\text{\tiny A}}_{t}}\!\!=\!\!\frac{\exp\!\left[\left(-i\nu t+2\lambda\int_{0}^{t}{\langle\hat{\sigma}_{z}\rangle}_{t}^{\text{\tiny A}}\text{d}s+\sqrt{\lambda}W(t)\right)\hat{\sigma}_{z}\right]\!\ket{\psi_{0}}}{\bra{\psi_{0}}\exp[(2\sqrt{\lambda}W(t)+4\lambda\int_{0}^{t}\langle\hat{\sigma}_{z}\rangle_{t}^{\text{\tiny A}}\text{d}s)\hat{\sigma}_{z}]\ket{\psi_{0}}^{1/2}},$ (14)
@@
 Adler et al.
-(2001) one can show that  [^8] See Appendix  E, where we show the collapse of the wavefunction in the nonlinear unraveling:
+(2001) one can show that  [^8] See Appendix  E, where we show the collapse of the wavefunction in the nonlinear unraveling.:
 $\lim_{t\rightarrow\infty}\Sigma^{\text{\tiny A}}_{t}(\hat{\sigma}_{z})=\lim_{t\rightarrow\infty}[\langle\hat{\sigma}_{z}^{2}\rangle^{\text{\tiny A}}_{t}-{\langle\hat{\sigma}_{z}\rangle^{\text{\tiny A}}_{t}}^{2}]=0,$ (15)
@@
 We now turn to the question of whether the nonlinear quantities introduced earlier can be operationally used to distinguish between different unravelings.
-From the analysis in the previous section, as well as existing literature Wiseman and Diósi (2001); Caiaffa et al.
-(2017), it is evident that these quantities differentiate between unravelings of the same master equation.
+From the analysis in the previous section, as well as existing literature Ghirardi et al.
+(1990b); Pearle (1997); Wiseman and Diósi (2001); Adler and Bassi (2007); Caiaffa et al.
+(2017), it is clear that these quantities differentiate between unravelings of the same master equation.
 However, this distinction has so far been purely mathematical.
@@
 This issue can be addressed at both the mathematical and operational levels.
-From a mathematical standpoint, it is well known that unravelings can be distinguished Wiseman and Diósi (2001); Caiaffa et al.
+From a mathematical standpoint, it is well known that unravelings can be distinguished Ghirardi et al.
+(1990b); Pearle (1997); Wiseman and Diósi (2001); Adler and Bassi (2007); Caiaffa et al.
 (2017), not least because the very notion of different unravelings implies some form of differentiation.